Please ignore the circled answers. I am checking my results.

rajoey is chose b. randen, whats the probability ios wegsboe of these. done a) o bX. 121 e 2266 ) .1056 ) none of these b).121 ) 2266 .1056 none of these b)) less than 1% c) 5% d) 0% e) none ofthese 32. If 1 joey is chosen at random, whats the probubility that it weis esasctly s00 mg a) 0 33. If 10 joeys are chosen at random, how likely is it that they average between 450 and 470 mg
Answer #1


Answer Date: 15/01/2019 31) If 1 Joey is chosen at random, the probability is that weighs between 450 and 470 mgis, First, coIf 1 Joey is chosen at random, the probability is that weighs between 450 and 470 mgis, First, compute the zscore then find pp (4505元 470-pC-1 .25s z -0.75) = p(25-075)-p(z -1.25) 0.2266-0.1057 = 0.1209~ 0.121 If 1 Joey is chosen at random, the proba500 500 40 0.00 From the standard nomal distribution table, the associated probability for the area to the left is shown beloThe mean ror sample mean 1s, 500 Standard deviation ofthe distribution of sample means is, 40 10 12.65 If 10 Joeys is chosen